# Fast-MCS: A Scalable Open-Source Tool to Find Minimal Cut Sets

### v1

@@ Abstract
A network is represented as a graph consisting of nodes and edges. A cut set for a source-destination pair in a network is a set of elements that, when failed, cause the source-destination pair to lose connectivity. A MCS (MCS) is a cut set that cannot be further reduced while maintaining its status as a cut set. MCS are crucial in identifying the critical elements in the network that have the most significant impact on failure. This work introduces Fast-MCS, an open-source, scalable tool for evaluating MCS in large, complex networks. Additionally, we compare the computation time of Fast-MCS with the state-of-the-art.
@@ I Introduction
The connectivity between any source and destination relies on how they are connected in the topology. When a set of elements in the network fails, disconnecting the source-destination pair, that set is called a cut set. A MCS is also a cut set that cannot be further reduced [1]. For example, consider the Germany_17 topology [2] shown in Fig. 1. Let us assume some traffic between the source Ulm and the destination Hamburg. For this source-destination pair, Ulm-Hamburg, (Frankfurt, Leipzig, Hannover) is a cut set because the failure of these nodes causes Ulm and Hamburg to be disconnected. However, the set can be further reduced to (Frankfurt, Leipzig) and still maintain its status as a cut set. Hence, (Frankfurt, Leipzig) is a MCS (MCS). Note that a single source-destination pair can have several MCS of varying sizes. For example, (Dortmund, Hannover, Leipzig) is another MCS of Ulm-Hamburg with size three. Moreover, a node can be part of several MCS. In this example, Leipzig is a part of both the MCS discussed.
@@
MCS are a popular tool used in modern networks for various purposes across various fields, including aerospace, nuclear power, transportation, electrical infrastructure, communication networks, etc. For example, MCS are used in Fault Tree Analysis [3, 4, 5, 6, 7], RBD (RBD) and availability evaluation [8, 9, 10], sovereignty metric evaluation [11, 12], resilience evaluation [13, 14], etc. Note that the network type does not change the concepts discussed in this work. Our Fast-MCS solution is applicable to any network or system that can be represented as a graph consisting of components as nodes and their interconnections as edges. However, this manuscript focuses on communication networks as the primary example.
Several nodes and edges exist in large communication networks. For a source and destination in the network, finding the MCS, i.e., finding the smallest combinations of nodes whose failure can cause loss of connection between that source and destination, gets computationally expensive as the size of the topology increases. This problem is noted to be NP-hard [15]. The aforementioned tools and evaluations depend on finding the MCS first. Therefore, it is essential to find a fast algorithm to evaluate MCS based on the topology information.
@@
    Sec. II investigates methods to evaluate MCS.
@@
    Sec. III discusses our novel method to evaluate find MCS, and its implementation as an open-source tool called Fast-MCS.
@@ II Background and Related Work
The most intuitive method to find MCS is a combinatorial searching algorithm, i.e., checking if each combination of nodes in the network could be a MCS. [16, 8, 17] propose extensions to the classical combinatorial searching algorithm to speed up the process. The authors in [18] extend this approach using Direct Partial Boolean Derivatives (DPBDs). They study the impact of changing a component’s state on the overall system availability, with positive and negative effects on path and cut sets, respectively. However, the aforementioned approaches do not scale well for large topologies because the number of possible node combinations scales as $\sum\limits_{i=1}^{|V|-1}\binom{|V|}{i}$.
@@
Other previous works also consider evaluating MCS. Most of these works discuss the evaluation of MCS based on MPS. A path set between a source and a destination is the complement of a cut set; i.e., a path set is a set of nodes whose availability connects the source and destination. A MPS is also a path set, but it cannot be further reduced. For example, let us assume some traffic between the source Ulm and the destination Hamburg, in the Germany_17 topology in Fig. 1. For this source-destination pair, {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hannover, Hamburg} is a path set because the availability of these nodes ensures a successful path between the source and the destination. However, this path set is not minimal because it can be further reduced to {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hamburg}. This new set cannot be reduced further, and hence is a MPS. Similar to MCS, a single source-destination pair can have several MPS. For example, {Ulm, Stuttgart, Nuernberg, Leipzig, Berlin, Hamburg} is another MPS. Additionally, a single node can be a part of multiple MPS.
@@
MPS and MCS are the prime implicants of monotone Boolean functions [19] for the success and the failure of a system, respectively. They are also referred to as the minimal satisfying and maximal non-satisfying vectors of a Boolean function [20]. Since a MPS is the logical opposite of a MCS, finding the logical complement of the MPS yields the MCS.
@@
In [21], the authors introduced the INMIN algorithm for inverting path sets and cut sets. However, the approach relies on exhaustive power-set enumeration and does not scale to large systems. Furthermore, the INMIN algorithm only evaluates the cut sets. It does not evaluate MCS. This would require further computation. The authors in [22, 23] introduced a recursive inversion technique based on De Morgan’s laws and Boolean absorption to convert MPS to MCS. Other works, such as [24, 25], use similar approaches to find MCS.
@@
In [26], the authors present their approach based on De Morgan’s rules and Boole-Shannon expansion to evaluate MCS from MPS. Their primary advantage is the ability to split the expression into smaller partitions for faster calculation and recursive absorption. Therefore, the computation required to obtain the minimum-sum-of-products expression of the MCS. Their approach is efficient for smaller networks. However, it does not scale for larger networks. This method is further explained in Sec. III-C.
@@
    Our method relies on logically complementing the MPS by using De Morgan’s laws and selects a single common element per step to enable scalability across large topologies.
@@
    Our method uses a faster, more memory-efficient set-theoretic approach to complement MPS, rather than the long Boolean expressions used in [26].
@@
The implementation of the combinatorial method [16, 8, 17] and the Boole-Shannon expansion [26] are explained in Sections III-A and III-C, respectively. Note that this manuscript considers only nodes to facilitate easier explanation. The concepts of MCS and MPS apply to links as well as edges. Furthermore, our tool Fast-MCS can include links in the MCS evaluation, if required.
Figure 2: Combinatorial procedure to evaluate MCS (MCS).
@@ III Methodology
This section discusses the implementation of the combinatorial method [16, 8, 17], the Boole-Shannon expansion [26] technique, and our Fast-MCS. For the Boole-Shannon expansion [26] and Fast-MCS, we use the same algorithm to find MPS.
III-A Combinatorial method to find MCS
The combinatorial method [16, 8, 17] to find the MCS for a source-destination pair is explained in Fig. 2. First, all the simple paths from the source to the destination are noted down as rows. Second, all nodes except the source and destination are treated as potential MCS and listed as columns (in blue). The table is populated as follows.
@@ III-A Combinatorial method to find MCS
Next, all possible two-node combinations are taken as columns (in red). The supersets of previously identified MCS are removed because these supersets can only be cutsets and not MCS. Then, the columns can be populated by using the logical ‘OR’ operation between the respective columns of the nodes in the combination. For example, the column for $v_{i},v_{j}$ is obtained by the logical ‘OR’ operation between the values in the columns for $v_{i}$ and $v_{j}$. After populating the red columns, the test for MCS is performed again by checking the existence of columns with all ones.
This procedure of populating columns for different node combinations and checking for MCS is repeated until no more columns are possible. A collection of all the all-ones columns provides all the MCS for that source-destination pair. For example, in Fig. 2, $(v_{1},v_{2})$ and $(v_{2},v_{3},v_{4})$ are MCS.
III-B Finding MPS (MPS)
We find the MPS based on an extension of the DFS (DFS) algorithm seen in Alg. 1 [9]. First, the branching starts at the source node. The newly explored node is added to the list of visited nodes in that branch. At every step of the branching procedure, two conditions are checked:
@@ III-B Finding MPS (MPS)
Hence, the DFS algorithm must be modified to terminate a branch if a new node is being visited that can already be visited from some other node in the visited list directly. This modification to the DFS algorithm, shown in line 10 of Alg. 1, speeds up the search for MPS. The branching procedure continues until all branches have been explored and terminated.
Algorithm 1 Find MPS (MPS) based on the modified DFS (DFS) algorithm.
@@
Figure 3: Example topology, source-destination pair, and its corresponding MPS (MPS).
@@
III-C Finding MCS (MCS) based on Boole-Shannon expansion
In this method, the system’s success, i.e., the source-destination pair’s successful connectivity, is represented as its MPS. Failure is represented by Boolean complementation [26]. A pivot element that appears most frequently is chosen to apply the Boole-Shannon expansion. Shannon expansion is applied as follows.
@@ III-C Finding MCS (MCS) based on Boole-Shannon expansion
where $x$ is the pivot element. The $S_{x=1}$ and $S_{x=0}$ expressions are recursively split using the Boole-Shannon expansion, until there are no unique pivot elements. The resulting final expression can be simplified and checked for absorption to obtain the MCS. For example, in the first step, considering the example shown in Fig. 3, for the source-destination pair S-T, the success function is,
@@
Then the expressions are complemented and considered for further expansion. The final expression is reduced by Boolean absorption to obtain the MCS.
III-D Finding MCS (MCS) using Fast-MCS
Algorithm 2 Fast-MCS’s MCS (MCS) computation using a Decision Tree
@@ III-D Finding MCS (MCS) using Fast-MCS
The methodology used in Fast-MCS is shown in Alg. 2. It is also explained using the example topology in Fig. 3. Consider the source-destination pair STS-T. The MPS for this source-destination pair are first evaluated based on the method discussed in Sec. III-B. For the example in Fig. 3, STS-T pair has four MPS: SABTS-A-B-T, SADEFTS-A-D-E-F-T, SCDEFTS-C-D-E-F-T, and SCDBTS-C-D-B-T. After removing the source and destination nodes from the MPS, the MPS are represented as sets: $\{A,B\}$, $\{A,D,E,F\}$, $\{C,D,E,F\}$, and $\{C,D,B\}$. This is the first step shown in the illustrated example in Fig. 4, and detailed in Line 2 of Alg. 2.
@@
The set on the left $\{A.B\}$ cannot be further split. However, the sets on the right can be further split with $C$ as the most repeated element, giving the third step in Fig. 4 based on the Split function in Line 18 of Alg. 2. No further branching is possible, and hence the obtained branches are leaf nodes. After the recursive splitting operations, the branches must be simplified before combining. For example, the leaf node $\{E,F\}$, and $\{B\}$ must be multiplied based on the EvaluateSelf and Multiply functions in Lines 57 and 66 of Alg. 2, respectively. The two sets are multiplied and checked for absorption to remove any subsets. In this case, they are evaluated to be $\{BE,BF\}$. After the leaf nodes are evaluated, they must be combined using the Combine function in Line 40 of Alg. 2. For example, the branch with $\{A,E,F\}$ is combined with the aforementioned evaluation of the leaf node. The combination operation computes the union of each pair of elements from the two sets and checks for absorption, removing any resulting subsets if present. The combination operation is recursively repeated to identify the MCS for the given source-destination pair. For the example topology in Fig. 3 and the source-destination pair STS-T, the MCS are $\{(A,D),(A,C),(B,D),(B,E),(B,F)\}$.
@@ III-E Differentiating Fast-MCS from Boole-Shannon expansion [26]
Though our Fast-MCS and the Boole-Shannon expansion used in [26] employ a recursive decomposition strategy, they differ fundamentally. The Boole-Shannon expansion is applied directly on the Boolean representation of the system’s success function (MPS in this work), manipulating logical expressions to derive MCS as the prime implicants of the complemented function. On the other hand, our Fast-MCS operates strictly on a set-theoretic representation of MPS and never constructs an explicit Boolean formula. Fast-MCS’s decision tree recursively partitions the sets into subsets that either contain or do not contain an element.
@@ III-F Workflow
Fig. 5 shows the workflow adopted in this work. Given an input topology, first, MPS are evaluated for each source-destination pair. Then, the approaches to find MCS- combinatorial method, Boole-Shannon expansion, and our Fast-MCS are evaluated. The times required to find MPS and MCS using the different approaches are measured and then compared.
@@ IV Results
Figure 6: Comparing the runtimes for Fast-MCS and the combinatorial approach [8, 17] (Sec. III-A). The blue and red bars correspond to finding MPS and MCS from MPS, respectively. Combined, the two bars together represent the total runtime for Fast-MCS. The orange bar corresponds to the combinatorial method. Fast-MCS outperforms the combinatorial method in every topology.

### v2

@@ Abstract
A network is represented as a graph consisting of nodes and edges. A cut set for a source-destination pair in a network is a set of elements that, when failed, cause the source-destination pair to lose connectivity. A Minimal Cut Set (MCS) is a cut set that cannot be further reduced while maintaining its status as a cut set. MCSs are crucial in identifying the critical elements in the network that have the most significant impact on failure. This work introduces Fast-MCS, an open-source, scalable tool for evaluating MCSs in large, complex networks. Additionally, we compare the computation time of Fast-MCS with the state-of-the-art.
@@ I Introduction
The connectivity between any source and destination relies on how they are connected in the topology. When a set of elements in the network fails, disconnecting the source-destination pair, that set is called a cut set. A MCS is also a cut set that cannot be further reduced [1]. For example, consider the Germany_17 topology [2] shown in Fig. 1. Let us assume some traffic between the source Ulm and the destination Hamburg. For this source-destination pair, Ulm-Hamburg, (Frankfurt, Leipzig, Hannover) is a cut set because the failure of these nodes causes Ulm and Hamburg to be disconnected. However, the set can be further reduced to (Frankfurt, Leipzig) and still maintain its status as a cut set. Hence, (Frankfurt, Leipzig) is a Minimal Cut Set (MCS). Note that a single source-destination pair can have several MCSs of varying sizes. For example, (Dortmund, Hannover, Leipzig) is another MCS of Ulm-Hamburg with size three. Moreover, a node can be part of several MCSs. In this example, Leipzig is a part of both the MCSs discussed.
@@
MCSs are a popular tool used in modern networks for various purposes across various fields, including aerospace, nuclear power, transportation, electrical infrastructure, communication networks, etc. For example, MCSs are used in Fault Tree Analysis [3, 4, 5, 6, 7], Reliability Block Diagram (RBD) and availability evaluation [8, 9, 10], sovereignty metric evaluation [11, 12], resilience evaluation [13, 14], etc. Note that the network type does not change the concepts discussed in this work. Our Fast-MCS solution is applicable to any network or system that can be represented as a graph consisting of components as nodes and their interconnections as edges. However, this manuscript focuses on communication networks as the primary example.
Several nodes and edges exist in large communication networks. For a source and destination in the network, finding the MCSs, i.e., finding the smallest combinations of nodes whose failure can cause loss of connection between that source and destination, gets computationally expensive as the size of the topology increases. This problem is noted to be NP-hard [15]. The aforementioned tools and evaluations depend on finding the MCSs first. Therefore, it is essential to find a fast algorithm to evaluate MCSs based on the topology information.
@@
    Sec. II investigates methods to evaluate MCSs.
@@
    Sec. III discusses our novel method to evaluate find MCSs, and its implementation as an open-source tool called Fast-MCS.
@@ II Background and Related Work
The most intuitive method to find MCSs is a combinatorial searching algorithm, i.e., checking if each combination of nodes in the network could be a MCS. [16, 8, 17] propose extensions to the classical combinatorial searching algorithm to speed up the process. The authors in [18] extend this approach using Direct Partial Boolean Derivatives (DPBDs). They study the impact of changing a component’s state on the overall system availability, with positive and negative effects on path and cut sets, respectively. However, the aforementioned approaches do not scale well for large topologies because the number of possible node combinations scales as $\sum\limits_{i=1}^{|V|-1}\binom{|V|}{i}$.
@@
Other previous works also consider evaluating MCSs. Most of these works discuss the evaluation of MCSs based on Minimal Path Sets. A path set between a source and a destination is the complement of a cut set; i.e., a path set is a set of nodes whose availability connects the source and destination. A MPS is also a path set, but it cannot be further reduced. For example, let us assume some traffic between the source Ulm and the destination Hamburg, in the Germany_17 topology in Fig. 1. For this source-destination pair, {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hannover, Hamburg} is a path set because the availability of these nodes ensures a successful path between the source and the destination. However, this path set is not minimal because it can be further reduced to {Ulm, Muenchen, Nuernberg, Leipzig, Berlin, Hamburg}. This new set cannot be reduced further, and hence is a MPS. Similar to MCSs, a single source-destination pair can have several MPSs. For example, {Ulm, Stuttgart, Nuernberg, Leipzig, Berlin, Hamburg} is another MPS. Additionally, a single node can be a part of multiple MPSs.
@@
MPSs and MCSs are the prime implicants of monotone Boolean functions [19] for the success and the failure of a system, respectively. They are also referred to as the minimal satisfying and maximal non-satisfying vectors of a Boolean function [20]. Since a MPS is the logical opposite of a MCS, finding the logical complement of the MPS yields the MCS.
@@
In [21], the authors introduced the INMIN algorithm for inverting path sets and cut sets. However, the approach relies on exhaustive power-set enumeration and does not scale to large systems. Furthermore, the INMIN algorithm only evaluates the cut sets. It does not evaluate MCSs. This would require further computation. The authors in [22, 23] introduced a recursive inversion technique based on De Morgan’s laws and Boolean absorption to convert MPSs to MCSs. Other works, such as [24, 25], use similar approaches to find MCSs.
@@
In [26], the authors present their approach based on De Morgan’s rules and Boole-Shannon expansion to evaluate MCSs from MPS. Their primary advantage is the ability to split the expression into smaller partitions for faster calculation and recursive absorption. Therefore, the computation required to obtain the minimum-sum-of-products expression of the MCSs. Their approach is efficient for smaller networks. However, it does not scale for larger networks. This method is further explained in Sec. III-C.
@@
    Our method relies on logically complementing the MPSs by using De Morgan’s laws and selects a single common element per step to enable scalability across large topologies.
@@
    Our method uses a faster, more memory-efficient set-theoretic approach to complement MPSs, rather than the long Boolean expressions used in [26].
@@
The implementation of the combinatorial method [16, 8, 17] and the Boole-Shannon expansion [26] are explained in Sections III-A and III-C, respectively. Note that this manuscript considers only nodes to facilitate easier explanation. The concepts of MCSs and MPSs apply to links as well as edges. Furthermore, our tool Fast-MCS can include links in the MCSs evaluation, if required.
Figure 2: Combinatorial procedure to evaluate Minimal Cut Sets (MCSs).
@@ III Methodology
This section discusses the implementation of the combinatorial method [16, 8, 17], the Boole-Shannon expansion [26] technique, and our Fast-MCS. For the Boole-Shannon expansion [26] and Fast-MCS, we use the same algorithm to find MPSs.
III-A Combinatorial method to find MCSs
The combinatorial method [16, 8, 17] to find the MCSs for a source-destination pair is explained in Fig. 2. First, all the simple paths from the source to the destination are noted down as rows. Second, all nodes except the source and destination are treated as potential MCSs and listed as columns (in blue). The table is populated as follows.
@@ III-A Combinatorial method to find MCSs
Next, all possible two-node combinations are taken as columns (in red). The supersets of previously identified MCSs are removed because these supersets can only be cutsets and not MCSs. Then, the columns can be populated by using the logical ‘OR’ operation between the respective columns of the nodes in the combination. For example, the column for $v_{i},v_{j}$ is obtained by the logical ‘OR’ operation between the values in the columns for $v_{i}$ and $v_{j}$. After populating the red columns, the test for MCSs is performed again by checking the existence of columns with all ones.
This procedure of populating columns for different node combinations and checking for MCSs is repeated until no more columns are possible. A collection of all the all-ones columns provides all the MCSs for that source-destination pair. For example, in Fig. 2, $(v_{1},v_{2})$ and $(v_{2},v_{3},v_{4})$ are MCSs.
III-B Finding Minimal Path Sets (MPSs)
We find the MPSs based on an extension of the Depth First Search (DFS) algorithm seen in Alg. 1 [9]. First, the branching starts at the source node. The newly explored node is added to the list of visited nodes in that branch. At every step of the branching procedure, two conditions are checked:
@@ III-B Finding Minimal Path Sets (MPSs)
Hence, the DFS algorithm must be modified to terminate a branch if a new node is being visited that can already be visited from some other node in the visited list directly. This modification to the DFS algorithm, shown in line 10 of Alg. 1, speeds up the search for MPSs. The branching procedure continues until all branches have been explored and terminated.
Algorithm 1 Find Minimal Path Sets (MPSs) based on the modified Depth First Search (DFS) algorithm.
@@
Figure 3: Example topology, source-destination pair, and its corresponding Minimal Path Sets (MPSs).
@@
III-C Finding Minimal Cut Sets (MCSs) based on Boole-Shannon expansion
In this method, the system’s success, i.e., the source-destination pair’s successful connectivity, is represented as its MPSs. Failure is represented by Boolean complementation [26]. A pivot element that appears most frequently is chosen to apply the Boole-Shannon expansion. Shannon expansion is applied as follows.
@@ III-C Finding Minimal Cut Sets (MCSs) based on Boole-Shannon expansion
where $x$ is the pivot element. The $S_{x=1}$ and $S_{x=0}$ expressions are recursively split using the Boole-Shannon expansion, until there are no unique pivot elements. The resulting final expression can be simplified and checked for absorption to obtain the MCSs. For example, in the first step, considering the example shown in Fig. 3, for the source-destination pair S-T, the success function is,
@@
Then the expressions are complemented and considered for further expansion. The final expression is reduced by Boolean absorption to obtain the MCSs.
III-D Finding Minimal Cut Sets (MCSs) using Fast-MCS
Algorithm 2 Fast-MCS’s Minimal Cut Set (MCS) computation using a Decision Tree
@@ III-D Finding Minimal Cut Sets (MCSs) using Fast-MCS
The methodology used in Fast-MCS is shown in Alg. 2. It is also explained using the example topology in Fig. 3. Consider the source-destination pair STS-T. The MPSs for this source-destination pair are first evaluated based on the method discussed in Sec. III-B. For the example in Fig. 3, STS-T pair has four MPSs: SABTS-A-B-T, SADEFTS-A-D-E-F-T, SCDEFTS-C-D-E-F-T, and SCDBTS-C-D-B-T. After removing the source and destination nodes from the MPSs, the MPSs are represented as sets: $\{A,B\}$, $\{A,D,E,F\}$, $\{C,D,E,F\}$, and $\{C,D,B\}$. This is the first step shown in the illustrated example in Fig. 4, and detailed in Line 2 of Alg. 2.
@@
The set on the left $\{A.B\}$ cannot be further split. However, the sets on the right can be further split with $C$ as the most repeated element, giving the third step in Fig. 4 based on the Split function in Line 18 of Alg. 2. No further branching is possible, and hence the obtained branches are leaf nodes. After the recursive splitting operations, the branches must be simplified before combining. For example, the leaf node $\{E,F\}$, and $\{B\}$ must be multiplied based on the EvaluateSelf and Multiply functions in Lines 57 and 66 of Alg. 2, respectively. The two sets are multiplied and checked for absorption to remove any subsets. In this case, they are evaluated to be $\{BE,BF\}$. After the leaf nodes are evaluated, they must be combined using the Combine function in Line 40 of Alg. 2. For example, the branch with $\{A,E,F\}$ is combined with the aforementioned evaluation of the leaf node. The combination operation computes the union of each pair of elements from the two sets and checks for absorption, removing any resulting subsets if present. The combination operation is recursively repeated to identify the MCSs for the given source-destination pair. For the example topology in Fig. 3 and the source-destination pair STS-T, the MCSs are $\{(A,D),(A,C),(B,D),(B,E),(B,F)\}$.
@@ III-E Differentiating Fast-MCS from Boole-Shannon expansion [26]
Though our Fast-MCS and the Boole-Shannon expansion used in [26] employ a recursive decomposition strategy, they differ fundamentally. The Boole-Shannon expansion is applied directly on the Boolean representation of the system’s success function (MPSs in this work), manipulating logical expressions to derive MCSs as the prime implicants of the complemented function. On the other hand, our Fast-MCS operates strictly on a set-theoretic representation of MPSs and never constructs an explicit Boolean formula. Fast-MCS’s decision tree recursively partitions the sets into subsets that either contain or do not contain an element.
@@ III-F Workflow
Fig. 5 shows the workflow adopted in this work. Given an input topology, first, MPSs are evaluated for each source-destination pair. Then, the approaches to find MCSs- combinatorial method, Boole-Shannon expansion, and our Fast-MCS are evaluated. The times required to find MPSs and MCSs using the different approaches are measured and then compared.
@@ IV Results
Figure 6: Comparing the runtimes for Fast-MCS and the combinatorial approach [8, 17] (Sec. III-A). The blue and red bars correspond to finding MPSs and MCSs from MPSs, respectively. Combined, the two bars together represent the total runtime for Fast-MCS. The orange bar corresponds to the combinatorial method. Fast-MCS outperforms the combinatorial method in every topology.